OpenStudy (anonymous):

SIMPLIFY AND STATE RESRICTIONS? \[\frac{2x^{2}-5x-3}{2x^{2}-11x+15} \times \frac{4^2-8x-5 }{4x^2+4x+1}\]

OpenStudy (anonymous):

Factor every element to get \[ \frac{2x^{2}-5x-3}{2x^{2}-11x+15} \times \frac{4x^2-8x-5 }{4x^2+4x+1}=\frac{((x-3) (2 x+1)) ((2 x-5) (2 x+1))}{((x-3) (2 x-5)) (2 x+1)^2} \] Cancel what you can cancel and state the restrictions

OpenStudy (anonymous):

When you cancel you get 1. Subject to the restrictions that none of the factors are zero. For example \[ x\ne 3\\ x\ne -\frac 1 2 \\ etc \]

OpenStudy (anonymous):

OH ! ... my goodness i factored wrong for 2x^2-11x+15... i put (2x+5)(x+3) kk so for restrictions... i would simplify first? then state restrictions?

OpenStudy (anonymous):

yes